OpenStudy (anonymous):

calculate the no of atoms in 0.5 mole atomes of nitrogen

OpenStudy (somy):

I think you need to use Avogadro constant- it is no. of atoms per mole so that means you just need to multiply Avogadro constant by number of moles Avogadro constant= 6.02*10^23 mole= 0.5mol so no. of atoms= (6.02*10^23)*(0.5) = 3.01*10^23 atoms

OpenStudy (anonymous):

thanks @somy..but wat is mole atoms....is it moles?

OpenStudy (ipwnbunnies):

Yeah, they're called moles, not mole atoms.

OpenStudy (somy):

@charutha i think you missed 'of' there, i mean its 0.5 mole of atoms of nitrogen :)